OpenStudy (anonymous):

Find values of c such that y=e^(cx) is a solution to the differential equation: y''-3y'+2y=0

OpenStudy (dumbcow):

set up characteristic equation ... which is quadratic where y'' is squared term , y' is "x" term and y is contant --> c^2 -3c +2 = 0 now solve for c

OpenStudy (anonymous):

Thanks, dumbcow. I had a feeling that the answer was found by using solving for a quadratic equation. So, how does find the zero's of the y function have anything to do with the y=e^x function?

OpenStudy (anonymous):

By y function I mean y'' -3y' +2y=0. Does it just so happen that the natural exponential function and this y'' -3y' +2y are related?

OpenStudy (dumbcow):

yeah it just so happens that the zeros are also the coefficients in y=e^cx here is a link that will help explain http://tutorial.math.lamar.edu/Classes/DE/RealRoots.aspx

OpenStudy (anonymous):

Thanks dumbcow, I appreciate all the help!

OpenStudy (dumbcow):

yw

OpenStudy (anonymous):

show the steps involved in working out the problem